How many revolutions will a car wheel of diameter 28 inches make as the car travels a distance of two miles?
Answer by rfer(16322)

5280*2=10560
10560*12=126720 ins
28*3.14=87.92 in circumfrence
126720/87.92=1441.31 revolutions